# MarS-FM: Generative Modeling of Molecular Dynamics via Markov State Models

### v2

@@ 4.1 A representative framework: Markov Space Flow Matching
Constructing MSMs. Specific versions of MarS-FM are instantiated by first building an MSM. To this aim, we use standards tools (Scherer et al., 2015; Hoffmann et al., 2021).
@@
The states $\{S_{i}\}$ are defined in a lower-dimensional space of collective variables. A common approach for dimensionality reduction is Time-lagged Independent Component Analysis (TICA) (Pérez-Hernández et al., 2013; Wu et al., 2017), which identifies directions $\mathbf{w}_{j}$ that maximize the autocorrelation of $\mathbf{w}_{j}^{\top}x(t)$ at lag time $\tau$. Alternatively, clustering can be applied directly to observables such as radius of gyration and fraction of secondary structures (Figure 2). Given states, the Markov chain transition matrix $\mathsf{T}$ is estimated at a lag time $\tau$. Two key observations are worth emphasizing. First, the MSM construction is performed once per dataset and remains fixed across all training trajectories, ensuring minimal computational overhead. Second, since MarS-FM learns to interpolate between MSM states beyond the transitions directly observed in MD, we can afford to select a larger $\tau$ than MD-Emus without sacrificing the data availability. We emphasize that MSMs are defined as pre-processing of training data; importantly, no MSM information is provided during inference. Additional details on how we construct MSMs are reported in Appendix B.
Representation. To generate protein conformations, MarS-FM adopts the same SE(3) representation as (Jumper et al., 2021; Jing et al., 2024b). Each residue $\ell$ of frame $x(t)$ is represented as
@@
where $q^{\ell}(t)\in\mathbb{R}^{4}$ is a unit quaternion describing a rotation and $r^{\ell}(t)\in\mathbb{R}^{3}$ is a translation, while $\theta_{k}^{\ell}(t)$ represent the $k$ torsion angles. Given an input $x(t)$, the target conformation $y$ is represented as an offset in roto-translational space. To model a vector field $v_{\theta}$ transporting a source distribution to the target distribution $p_{\mathsf{T}}(\cdot|x(t))$, we build on the MDGen architecture (Jing et al., 2024b). Specifically, $v_{\theta}$ leverages DiT-style blocks (Peebles & Xie, 2023) and is conditioned on protein sequence $a$ and the current conformation $x(t)$ via IPA layers (Jumper et al., 2021)—see Appendix A for a complete description. For notational simplicity, we omit the explicit conditioning on the protein sequence $a$.
Training objective. We train MarS-FM using Flow Matching, which learns a time-dependent vector field $v_{\theta}$ to transport a source distribution $p_{0}={\mathcal{N}}(0,1)$ to the target distribution $p_{1}(\cdot|x(t))=p_{\mathsf{T}}(\cdot|x(t))$. Concretely, given an input frame $x(t)\in S_{i}$, we first draw a state $S_{j}$ according to the probability distribution $j\mapsto\mathsf{T}_{ij}$ and then sample $x_{1}\in S_{j}$ uniformly within such a state. Given sample pairs $x_{0}\sim p_{0}$ and $x_{1}\sim p_{1}$, we define interpolations $0,1]\ni s\mapsto x_{s}$ connecting $x_{0}$ to $x_{1}$. We optimize $v_{\theta}$ by minimizing velocity mismatch along conditional paths $s\mapsto p_{s}(\cdot|x_{0},x_{1})$ based on these interpolations (Lipman et al., 2022; Liu et al., 2022; Albergo & Vanden-Eijnden, 2022). That is, we consider the training loss
@@
where $\dot{x}_{s}$ is the velocity of the conditional path. Crucially, we have emphasized how we select each state $S_{i}$ uniformly, before conditioning on $x(t)$. This state-based sampling ensures that rare states are encountered more frequently during training. In contrast, standard MD-Emus draw frames uniformly from the trajectory, which often biases training toward common intra-state transitions. A similar strategy was used by Costa et al. (2024), though their model still approximates a fixed lag time transition. Further implementation and batching details are provided in Appendix A.
@@
2. MarS-FM $\oplus$ MD-Emu. We combine MarS-FM with MDGen (Jing et al., 2024b). We first sample conformations using MarS-FM and then use MDGen to generate shorter trajectories from each of these points, separately. This approach mirrors the MSM paradigm, where statistics are inferred by initiating short simulations from different states. We introduce such a hybrid scheme, as in certain workflows temporal fidelity may also be relevant (De Vivo et al., 2016).
@@ 5 Experiments
A key condition for a generative model to replace MD-sampling, is that measurements of observables under the generated samples match those under MD. For this reason, in our experiments, we compare distribution of observables computed along MD trajectories and those computed along conformations generated by the underlying model. We rely on observables used in previous generative works (Jing et al., 2024a, b) as well as additional ones reported for analyzing the MD-Cath dataset (Mirarchi et al., 2024).
Variants of MarS-FM. We first construct an MSM—dependent on the dataset but invariant across trajectories, as described below—and then train MarS-FM to sample from $p_{\mathsf{T}}(\cdot|x(t))$. Yet no information about MSMs is provided during inference. As explained in §4, during sampling, we consider two variants: MarS-FM $\oplus$ MDGen and MarS-FM.
Baselines. In our experiments, we take MDGen (Jing et al., 2024b) as the main representative of the MD-Emu class. Primarily, this is due to MDGen being the most versatile variant of MD-Emus due to the choice of the window $K$ of consecutive transitions at resolution $\tau$. For larger systems, we additionally compare against BioEmu (Lewis et al., 2024), noting specific caveats discussed in Appendix B.7. Finally, we point out that other methods are either not scalable for MD-Cath domain (Klein et al., 2023; Schreiner et al., 2023) or have only been validated within the training distribution (Costa et al., 2024) (see Appendix B.7 for further discussion).
@@ 5.1 Preliminary investigation: Tetrapeptides
We first adopt the dataset of tetrapeptides that provided the main evaluation for MDGen (Jing et al., 2024b). This consists of $\sim$3000 trajectories in training, 100 in validation and 100 in test. Each trajectory is simulated in explicit solvent for 100 ns and conformations are saved every 10 ps. That is, for any tetrapeptide in the test set, we have a held-out MD distribution of 104 frames. Accordingly, we generate distributions of 104 frames using both MD-Emus (MDGen) and MSM-Emus (MarS-FM).
MSM construction. When training MarS-FM, we need to define MSMs over the MD trajectories. As this dataset contains exhaustive simulations relative to the size of the systems, we rely on TICA (Pérez-Hernández et al., 2013), selecting the minimal number of TICA coordinates whose cumulative kinetic variance exceeds 95%. We then apply k-means clustering to these coordinates, obtaining 100 microstates, which we further group into 10 metastable states using the PCCA+ spectral clustering method (Röblitz & Weber, 2013). This procedure yields the final MSMs constructed with a lag time of 100 ps. Further details on MSM construction are provided in Appendix B.
Evaluation and metrics. During sampling, we report the Jensen-Shannon Divergence (JSD) over different observables. Namely, we follow the same metrics as in MDGen Jing et al. (2024b): (i) Torsional angles (of both backbone and sidechains); (ii) Projection onto the slowest modes as identified by TICA; (iii) Equilibrium distributions induced by MSMs built over the test peptides. Inspired by Lewis et al. (2024), we also report the Macrostate MAE (mMAE), which measures the mean absolute error between model-generated and ground-truth MD free energies across metastable states. The mMAE provides a direct measure of the model’s accuracy in reproducing metastable state distributions. Further computational details and definitions are presented in Appendix B.
@@
Torsions ($b$) Torsions (sc) Torsions (all) TICA-0 TICA-0,1 joint MSM states Macrostate MAE MD (Oracle) 0.10 0.06 0.08 0.20 0.27 0.21 MDGen-1000 0.13 0.09 0.11 0.23 0.32 0.23 1.13 MDGen-200 0.14 0.10 0.12 0.24 0.33 0.27 1.12 MarS-FM $\oplus$ MDGen-200 (ours) 0.12 0.09 0.10 0.21 0.30 0.23 0.83 MarS-FM (ours) 0.11 0.09 0.10 0.21 0.29 0.22 0.63
@@ 5.2 MD-Cath: Proteins with sequence dissimilarity and large domain motions
Our main evaluation is based on the MD-Cath dataset from Mirarchi et al. (2024). This includes $5398$ domains up to $500$ residues (average number of residues $137$), simulated at 55 different temperatures, spanning 55 random replicas per setting. The average length of each trajectory is $464$ ns and frames are saved at 11 ns intervals. The dataset is already built to ensure protein diversity by considering non-homologous domains at the $S20$ ($20\%$) homology level—further details are reported in Appendix B.
MSM construction. MSMs are constructed independently for each domain using k-means clustering directly applied to normalized observables identified as critical by the original dataset analysis (Mirarchi et al., 2024). Specifically, we consider two features: (i) the radius of gyration and (ii) the fraction of residues in $\alpha$-helical and $\beta$-sheet secondary structures—additional details are reported in Appendix B. We cluster these normalized features into 10 metastable states, providing a physically meaningful partitioning of the conformational space. This yields the final MSMs constructed with a lag time of 50 ns.
Large domain motions. To validate the ability of MarS-FM to generate large conformational changes, our main evaluation is performed over the highest-temperature replica ($450$ K), in which proteins exhibit unfolding. In fact, proteins are quite stable at lower temperatures over the simulated timescales, meaning that often no significant conformational change has occurred—as reported in Mirarchi et al. (2024)—which would result in an easier generative task. Nonetheless, for a more complete evaluation, we also train MarS-FM on lowest-temperature replica and sample conformations in this regime–results are reported in Appendix C.
Assessing generalization. First, we consider a random training-validation-test split partitioning the total number of domains according to $80\%$-$\%-1010\%$. We then follow the strategy in Lewis et al. (2024) and use mmseqs2 (Steinegger & Söding, 2017) with highest sensitivity setting to filter out any test protein sharing more than $20\%$ sequence similarity with any training protein (note that Lewis et al. (2024) use a more forgiving $40\%$ threshold). As a result, our test set consists of $495$ domains that are meaningfully diverse from those whose trajectories have been used during training.
@@
Metrics and Observables. For evaluation, we focus on observables studied in the original dataset analysis of Mirarchi et al. (2024), namely the radius of gyration and fractions of residues in secondary structures ($\alpha$-helices and $\beta$-sheets). We measure the forward KL divergence to quantify the model’s exploration capabilities and the Jensen-Shannon Divergence (JSD) for general distribution alignment. Inspired by Jing et al. (2024a), we also quantify ensemble flexibility using the Pearson correlation ($r$) computed on pairwise backbone Root Mean Square Deviation (RMSD), global Root Mean Square Fluctuation (RMSF), and per-target RMSF. We additionally evaluate folding free energies via the mean absolute error ($\Delta G_{\text{fold}}$ MAE). Lastly, we calculate the JSD between reconstructed MSMs and reference MSMs, providing a direct measure of the ability to explore states of domains dissimilar from those observed in training. Additional details on metrics are provided in Appendix B.
@@
Extended baselines. We benchmark different variants of MDGen to highlight how advantages of our framework cannot be simply replicated by changing the lag time or altering the sampling approach. As such, we train MDGen with window size $K=20$ and $K=100$, corresponding to a total lag time of 2020 ns and $100$ ns, respectively. Additionally, to demonstrate that the improvements of MarS-FM cannot be simply ascribed to a reduction of autoregressive calls, we also evaluate MDGen in parallel, meaning that we sample conformations only conditioned on the input frame. Finally, we also report the performance of one MD replica against 4 held-out ones (and average over all possible 5 combinations) to be treated as oracle performance.
@@
Sampling speed. A key requirement for generative models to serve as practical replacements for MD is that they offer substantial computational speed-ups.
@@
Table 4 summarizes the wall-clock time required to generate the equivalent of 500 ns of trajectory (500 conformations) for a 159-residue protein (domain 4dhkB00) on an NVIDIA H100. For a fair comparison, we report MD in implicit solvent, which is substantially more efficient than explicit-solvent simulations. For such a system, implicit MD runs at $\approx 2400$ ns/day on the same hardware for a 159-residue protein (OpenMM Team, 2025). Under these conditions, MarS-FM provides a computational speed-up of $600\times$.

### v3

@@ 4.1 A representative framework: Markov Space Flow Matching
Constructing MSMs. Specific versions of MarS-FM are instantiated by first building an MSM. To this aim, we use standard tools (scherer2015pyemma; hoffmann2021deeptime).
@@
The states $\{S_{i}\}$ are defined in a lower-dimensional space of collective variables. A common approach for dimensionality reduction is Time-lagged Independent Component Analysis (TICA) (perez2013identification; wu2017variational), which identifies directions $\mathbf{w}_{j}$ that maximize the autocorrelation of $\mathbf{w}_{j}^{\top}x(t)$ at lag time $\tau$. Alternatively, clustering can be applied directly to observables such as radius of gyration and fraction of secondary structures (Figure 2). Given states, the Markov chain transition matrix $\mathsf{T}$ is estimated at a lag time $\tau$.
Two key observations are worth emphasizing. First, MSM construction is carried out once as an offline pre-processing step: for each dataset we fix the MSM hyperparameters and then build a separate MSM for every training domain from its MD trajectories. These per-domain state spaces and transition matrices remain fixed throughout training, ensuring minimal computational overhead. Second, since MarS-FM learns to interpolate between MSM states beyond the transitions directly observed in MD, we can afford to select a larger $\tau$ than MD-Emus without sacrificing the data availability. The lag time still sets the coarse-grained timescale of the underlying Markov chain, but the number and diversity of effective training transitions is no longer tied to how often frames are saved in the MD trajectories: any pair of frames assigned to successive MSM states can be used to define a training pair.
We emphasize that MSMs are defined as pre-processing of training data; importantly, no MSM information is provided during inference. Additional details on how we construct MSMs are reported in Appendix B.
Representation. To generate protein conformations, MarS-FM adopts the same SE(3) representation as (jumper2021highly; jing24generative). Each residue $\ell$ of frame $x(t)$ is represented as
@@
where $q^{\ell}(t)\in\mathbb{R}^{4}$ is a unit quaternion describing a rotation and $r^{\ell}(t)\in\mathbb{R}^{3}$ is a translation, while $\theta_{k}^{\ell}(t)$ represent the $k$ torsion angles. Given an input $x(t)$, the target conformation $y$ is represented as an offset in roto-translational space. To model a vector field $v_{\theta}$ transporting a source distribution to the target distribution $p_{\mathsf{T}}(\cdot|x(t))$, we build on the MDGen architecture (jing24generative). Specifically, $v_{\theta}$ leverages DiT-style blocks (peebles2023scalable) and is conditioned on protein sequence $a$ and the current conformation $x(t)$ via IPA layers (jumper2021highly)—see Appendix A for a complete description. For notational simplicity, we omit the explicit conditioning on the protein sequence $a$.
Training objective. We train MarS-FM using Flow Matching, which learns a time-dependent vector field $v_{\theta}$ to transport a source distribution $p_{0}={\mathcal{N}}(0,1)$ to the target distribution $p_{1}(\cdot|x(t))=p_{\mathsf{T}}(\cdot|x(t))$. Concretely, given an input frame $x(t)\in S_{i}$, we first draw a destination state $S_{j}$ according to the MSM transition probabilities $j\mapsto\mathsf{T}_{ij}$ and then choose a target MD frame among those assigned to $S_{j}$. In practice we sample uniformly from the finite set of MD frames in $S_{j}$. We denote the resulting target frame by $x_{1}$.
Independently, we draw a noise sample $x_{0}\sim p_{0}$ and define interpolations $0,1]\ni s\mapsto x_{s}$ connecting $x_{0}$ to $x_{1}$ using the standard schedule of Flow Matching (see Appendix A.2 for the explicit formula for $x_{s}$ and its analytic velocity $\dot{x}_{s}$). We optimize $v_{\theta}$ by minimizing the velocity mismatch along the conditional paths $s\mapsto p_{s}(\cdot\mid x_{0},x_{1})$ induced by these interpolations (lipman_flow_2022; liu2022flow; albergo2022building). That is, we consider the training loss
@@
where $\dot{x}_{s}$ is the velocity of the conditional path. Crucially, we first select each state $S_{i}$ uniformly (see Appendix A.1) before conditioning on a specific frame $x(t)\in S_{i}$. This state-based sampling ensures that rare states are encountered more frequently during training. In contrast, standard MD-Emus draw frames uniformly from the trajectory, which often biases training toward common intra-state transitions. A similar strategy was used by costa2024equijump, though their model still approximates a fixed lag time transition. Further implementation and batching details are provided in Appendix A.
Compared to MDGen, which always takes $x_{1}$ to be a future frame $x(t+K\tau)$ from the same trajectory, MarS-FM samples $x_{1}$ via the MSM transition kernel $\mathsf{T}$, so that the same Flow Matching objective is applied to a more diverse, state-conditioned set of training pairs. In particular, the model is not trained solely on transitions of the form $x(t+K\tau)$: the MSM provides transitions between all pairs of frames assigned to successive metastable states. Because each metastable state aggregates frames from different replicas and from multiple visits to the same basin within a trajectory, this construction exposes MARS-FM to transitions that span replicas and multiple barrier-crossing events, even if only a few such events occur in any individual time series. The lag time used to define the MSM still determines the coarse-grained timescale of these moves, but the effective training set is no longer restricted to the specific $x(t)\to x(t+\tau)$ pairs obtained at a fixed frame-saving interval. While these transitions still implicitly correspond to a characteristic timescale set by the MSM lag time, the number and diversity of transitions available for training are no longer constrained by the temporal resolution used to slice the trajectories.
@@
2. MarS-FM $\oplus$ MD-Emu. We combine MarS-FM with MDGen (jing24generative). We first sample conformations using MarS-FM and then use MDGen to generate shorter trajectories from each of these points, separately. This approach mirrors the MSM paradigm, where statistics are inferred by initiating short simulations from different states. We introduce such a hybrid scheme, as in certain workflows temporal fidelity may also be relevant (de2016role).
@@ 5 Experiments
A key condition for a generative model to replace MD-sampling, is that measurements of observables under the generated samples match those under MD. For this reason, in our experiments, we compare distribution of observables computed along MD trajectories and those computed along conformations generated by the underlying model. We rely on observables used in previous generative works (jing2024alphafold; jing24generative) as well as additional ones reported for analyzing the MD-Cath dataset (mirarchi2024mdcath).
Variants of MarS-FM. We first construct per-domain MSMs with dataset-specific hyperparameters—and then train MarS-FM to sample from $p_{\mathsf{T}}(\cdot|x(t))$. Yet no information about MSMs is provided during inference. As explained in §4, during sampling, we consider two variants: MarS-FM $\oplus$ MDGen and MarS-FM.
Baselines. In our experiments, we take MDGen (jing24generative) as the main representative of the MD-Emu class. Primarily, this is due to MDGen being the most versatile variant of MD-Emus due to the choice of the window $K$ of consecutive transitions at resolution $\tau$. For larger systems, we additionally compare against BioEmu (lewis2024scalable), noting specific caveats discussed in Appendix B.7. Finally, we point out that other methods are either not scalable for MD-Cath domain (klein2023timewarp; schreiner2023implicit) or have only been validated within the training distribution (costa2024equijump) (see Appendix B.7 for further discussion).
@@ 5.1 Preliminary investigation: Tetrapeptides
We first adopt the dataset of tetrapeptides that provided the main evaluation for MDGen (jing24generative). This consists of $\sim$3000 trajectories in training, 100 in validation and 100 in test. Each trajectory is simulated in explicit solvent for 100 ns and conformations are saved every 10 ps. That is, for any tetrapeptide in the test set, we have a held-out MD distribution of 104 frames. Accordingly, we generate distributions of 104 frames using both MD-Emus (MDGen) and MSM-Emus (MarS-FM).
MSM construction. When training MarS-FM, we need to define MSMs over the MD trajectories. As this dataset contains exhaustive simulations relative to the size of the systems, we rely on TICA (perez2013identification), selecting the minimal number of TICA coordinates whose cumulative kinetic variance exceeds 95%. We then apply k-means clustering to these coordinates, obtaining 100 microstates, which we further group into 10 metastable states using the PCCA+ spectral clustering method (roblitz2013fuzzy). This procedure yields the final MSMs constructed with a lag time of 2 ns. Further details on MSM construction are provided in Appendix B.
Torsions (bb) Torsions (sc) Torsions (all) TICA-0 TICA-0,1 joint MSM states Macrostate MAE MD (Oracle) 0.10 0.06 0.08 0.20 0.27 0.21 MDGen-1000 0.13 0.09 0.11 0.23 0.32 0.23 1.13 MDGen-200 0.14 0.10 0.12 0.24 0.33 0.27 1.12 MarS-FM $\oplus$ MDGen-200 (ours) 0.12 0.09 0.10 0.21 0.30 0.23 0.83 MarS-FM (ours) 0.11 0.09 0.10 0.21 0.29 0.22 0.63
Evaluation and metrics. During sampling, we report the Jensen-Shannon Divergence (JSD) over different observables. Namely, we follow the same metrics as in MDGen jing24generative: (i) Torsional angles (of both backbone and sidechains); (ii) Projection onto the slowest modes as identified by TICA; (iii) Equilibrium distributions induced by MSMs built over the test peptides. Inspired by lewis2024scalable, we also report the Macrostate MAE (mMAE), which measures the mean absolute error between model-generated and ground-truth MD free energies across metastable states. The mMAE provides a direct measure of the model’s accuracy in reproducing metastable state distributions. Further computational details and definitions are presented in Appendix B.
@@ 5.2 MD-Cath: Proteins with sequence dissimilarity and large domain motions
Our main evaluation is based on the MD-Cath dataset from mirarchi2024mdcath. This includes $5398$ domains up to $500$ residues (average number of residues $137$), simulated at $5$ different temperatures, spanning $5$ random replicas per setting. The average length of each trajectory is $464$ ns and frames are saved at $1$ ns intervals. The dataset is already built to ensure protein diversity by considering non-homologous domains at the $S20$ ($20\%$) homology level—further details are reported in Appendix B.
MSM construction. MSMs are constructed independently for each domain using k-means clustering directly applied to normalized observables identified as critical by the original dataset analysis (mirarchi2024mdcath). Specifically, we consider two features: (i) the radius of gyration and (ii) the fraction of residues in $\alpha$-helical and $\beta$-sheet secondary structures—additional details are reported in Appendix B. We cluster these normalized features into 10 metastable states, providing a physically meaningful partitioning of the conformational space. This yields the final MSMs constructed with a lag time of 50 ns.
@@
Large domain motions. To validate the ability of MarS-FM to generate large conformational changes, our main evaluation is performed over the highest-temperature replica ($450$ K), in which proteins exhibit unfolding. In fact, proteins are quite stable at lower temperatures over the simulated timescales, meaning that often no significant conformational change has occurred—as reported in mirarchi2024mdcath—which would result in an easier generative task. Nonetheless, for a more complete evaluation, we also train MarS-FM on lowest-temperature replica and sample conformations in this regime–results are reported in Appendix C.
Assessing generalization. First, we consider a random training-validation-test split partitioning the total number of domains according to $80\%$-$\%-1010\%$. We then follow the strategy in lewis2024scalable and use mmseqs2 (steinegger2017mmseqs2) with highest sensitivity setting to filter out any test protein sharing more than $20\%$ sequence similarity with any training protein (note that lewis2024scalable use a more forgiving $40\%$ threshold). As a result, our test set consists of $495$ domains that are meaningfully diverse from those whose trajectories have been used during training.
@@
Metrics and Observables. For evaluation, we focus on observables studied in the original dataset analysis of mirarchi2024mdcath, namely the radius of gyration and fractions of residues in secondary structures ($\alpha$-helices and $\beta$-sheets). We measure the forward KL divergence to quantify the model’s exploration capabilities and the Jensen-Shannon Divergence (JSD) for general distribution alignment. Inspired by jing2024alphafold, we also quantify ensemble flexibility using the Pearson correlation ($r$) computed on pairwise backbone Root Mean Square Deviation (RMSD), global Root Mean Square Fluctuation (RMSF), and per-target RMSF. We additionally evaluate folding free energies via the mean absolute error ($\Delta G_{\text{fold}}$ MAE). Lastly, we calculate the JSD between reconstructed MSMs and reference MSMs, providing a direct measure of the ability to explore states of domains dissimilar from those observed in training. Additional details on metrics are provided in Appendix B.
Extended baselines. We benchmark different variants of MDGen to highlight how advantages of our framework cannot be simply replicated by changing the lag time or altering the sampling approach. As such, we train MDGen with window size $K=20$ and $K=100$, corresponding to a total lag time of $20$ ns and $100$ ns, respectively. Additionally, to demonstrate that the improvements of MarS-FM cannot be simply ascribed to a reduction of autoregressive calls, we also evaluate MDGen in parallel, meaning that we sample conformations only conditioned on the input frame. Finally, we also report the performance of one MD replica against 4 held-out ones (and average over all possible 5 combinations) to be treated as oracle performance.
@@
Sampling speed. A key requirement for generative models to serve as practical replacements for MD is that they offer substantial computational speed-ups. Table 4 summarizes the wall-clock time required to generate the equivalent of 500 ns of trajectory (500 conformations) for a 159-residue protein (domain 4dhkB00) on an NVIDIA H100. For a fair comparison, we report MD in implicit solvent, which is substantially more efficient than explicit-solvent simulations. For such a system, implicit MD runs at $\approx 2400$ ns/day on the same hardware for a 159-residue protein (openmm_benchmarks). Under these conditions, MarS-FM provides a computational speed-up of $600\times$.